### v1

@@ Abstract
The integration of Large Language Models (LLMs) like GPT-4 into traditional Natural Language Processing (NLP) tasks has opened new avenues for enhancing model performance while reducing the reliance on extensive human annotations. This paper presents a novel approach that leverages the Chain of Thought (CoT) prompting technique to distill knowledge from GPT-4, subsequently applying it to improve the efficiency and effectiveness of a smaller model, BERT, on Named Entity Recognition (NER) tasks. Our method involves a two-phase training process: initially employing GPT-4 annotated data for pre-training and then refining the model with a combination of distilled and original human-annotated data. The results demonstrate that our mixed-training strategy significantly outperforms models trained solely on human annotations, achieving superior F1-scores and showcasing a cost-effective solution for resource-limited or closed-network settings. The study also discusses the challenges encountered, such as LLM output variability and the tendency towards hallucinations, proposing future work directions to enhance prompt design and annotation selection. Our findings indicate a promising synergy between LLM insights and traditional NLP techniques, paving the way for more accessible and robust NLP applications.
@@ Introduction
Addressing these challenges, our research proposes a comprehensive solution utilizing knowledge distillation techniques. By leveraging LLMs for data annotation and incorporating these annotations in the training of smaller models like BERT, we enhance the performance of these models on traditional NLP tasks. Our approach mitigates the limitations of LLMs through the use of Chain of Thought (CoT) prompting to refine reasoning processes and generate more accurate annotations. Furthermore, we introduce an efficient prompt design that effectively translates LLM outputs into usable formats for sequence labeling tasks, overcoming issues of entity overlap and discontinuation common in many domains. Through a series of experiments utilizing datasets such as CONLL and Xtreme, we demonstrate the effectiveness of our mixed training strategy, which combines distilled LLM annotations with manually annotated data, achieving superior performance over models trained solely on human-annotated datasets.
@@
The integration of Large Language Models (LLMs) into Named Entity Recognition (NER) tasks has seen varied approaches and methodologies, demonstrating the evolving landscape of NLP research in leveraging advanced AI models for enhancing traditional tasks. This section outlines related work that has contributed to the development of methods and strategies in applying LLMs and other advanced techniques to NER, highlighting the diverse efforts to address challenges such as data scarcity, model efficiency, and cross-lingual applications.
@@
Cross-lingual NER remains a significant challenge due to the scarcity of annotated training data, especially in low-resource languages. To address this, [1] introduces a Cross-lingual Entity Projection framework (CROP) that leverages a multilingual labeled sequence translation model for zero-shot cross-lingual NER, demonstrating substantial improvements over previous approaches. The gap between the sequence labeling nature of NER tasks and the text-generation orientation of LLMs is explored in [2], where the GPT-NER model is proposed. By transforming the NER task into a text generation format, GPT-NER addresses the hallucination issue common in LLMs, showing promise in low-resource and few-shot settings. PromptNER [3] presents a novel few-shot and cross-domain NER algorithm that utilizes prompt-based heuristics with LLMs, requiring only a few-shot examples and entity type definitions to achieve state-of-the-art performance on several benchmarks. This underscores the potential of prompt-based methods in NER tasks. [4] explores targeted distillation with mission-focused instruction tuning to train student models for NER tasks, demonstrating that these models can significantly outperform general instruction-tuned models across a broad application class, including open information extraction. The Multi-view Contrastive Learning approach for Cross-lingual NER (MCL-NER) proposed in [5] tackles the challenge of semantic and token-level representation alignment across languages, achieving new state-of-the-art performance on the XTREME benchmark for 40 languages. Inspired by the Chain of Thought methodology, [6] leverages the step-by-step problem-solving capabilities of LLMs for Biomedical NER, incorporating external knowledge to enhance entity type determination, showcasing significant improvements in a domain-specific application. [8] focuses on zero-shot information extraction in NER, adapting reasoning strategies tailored for NER tasks with LLMs. The introduction of syntactic augmentation and a two-stage majority voting strategy demonstrates remarkable improvements across multiple benchmarks. Lastly, [9] addresses the issue of applying pre-trained transformer models to long documents for NER tasks. By generating a synthetic context retrieval training dataset and training a neural context retriever based on a BERT model, this work shows that it is possible to outperform unsupervised retrieval baselines, marking an advancement in handling long-document NER tasks.
These works collectively underscore the dynamic interplay between LLMs, traditional NLP techniques, and innovative methodologies in advancing the field of NER. Our research builds upon these foundations, aiming to further leverage the capabilities of LLMs to enhance NER model training and performance, particularly through the integration of Chain of Thought prompting and knowledge distillation techniques.
@@
This study investigates the potential of enhancing traditional Natural Language Processing (NLP) tasks through the integration of Large Language Models (LLMs), specifically using GPT-4 for annotation and BERT for model training. Focusing on the Named Entity Recognition (NER) task within the widely recognized CONLL2003 benchmark, our approach is divided into two main phases. In the first phase, GPT-4 is employed to automatically annotate a selected subset of the dataset without any fine-tuning, aiming to assess the feasibility and quality of LLM-generated annotations. The second phase examines the impact of these LLM annotations on the performance of the BERT model, which is trained using a mix of original and LLM-annotated data. This comparative analysis seeks to uncover the benefits and limitations of leveraging LLM annotations in the context of traditional NLP frameworks, with the ultimate goal of improving efficiency and accuracy in language processing tasks.
Phase One: Data Annotation with LLM
@@
For our annotation process, we leveraged the GPT-4 turbo API to meticulously generate named entity annotations for the selected CONLL2003 data samples. This process involved employing two distinct prompting methods (detailed in subsequent sections), which facilitated the model’s understanding and execution of the annotation task. While GPT-4’s outputs were predominantly accurate and aligned with the task requirements, a minimal amount of manual post-processing was necessary to rectify occasional irregularities in the generated annotations, ensuring the highest quality of data for our study.
@@
We employed a few-shot prompting approach, shown in Figure 1(left), with GPT-4 to annotate a random sample of 1000 sentences from the CONLL2003 dataset, serving as a baseline for comparison. Few-shot learning, a method where the model learns to perform tasks from a limited number of examples, was utilized to guide GPT-4 in accurately identifying and classifying named entities into four categories: LOC (Location), MISC (Miscellaneous), ORG (Organization), and PER (Person).
@@
Alongside the standard prompting approach, we implemented the Chain of Thought (CoT) prompting method combined with few-shot learning to annotate the same 1000 randomly sampled sentences from the CONLL2003 dataset using GPT-4, which is shown in Figure 1(right), . The CoT methodology is designed to mimic human-like reasoning by guiding the model through a stepwise thought process before arriving at a conclusion. This approach not only aims to improve the accuracy of the annotations by making the model’s thought process explicit but also seeks to provide insights into the model’s reasoning, thereby enhancing the interpretability of the annotations.
@@
Figure 1: The Standard Prompting Template (left) and CoT prompting Template (right)
@@
In the subsequent phase of our research, we turned to the renowned BERT base model, specifically the uncased variant (bert-base-uncased), for fine-tuning with the data annotated by GPT-4. BERT, or Bidirectional Encoder Representations from Transformers, has been a transformative force in NLP, known for achieving state-of-the-art results across various tasks. Its architecture employs 12 layers of bidirectional transformers, making it adept at understanding the nuances of language context.
BERT’s uncased model, which was pre-trained on a large corpus, including the entirety of the English Wikipedia and the BookCorpus dataset, using a self-supervised learning approach. This approach involved two pre-training tasks: Masked Language Modeling (MLM) and Next Sentence Prediction (NSP), enabling the model to learn a deep bidirectional representation of the English language.
For our experiment, BERT’s bidirectional training framework makes it an ideal candidate to further distill and refine the knowledge transferred from LLMs like GPT-4. Its capacity to handle context effectively and its relatively compact architecture make it well-suited for scenarios with limited computational resources or closed-network environments, where leveraging the vast knowledge distilled from larger models like GPT-4 could be significantly beneficial.
The model we employed was structured to perform token classification, a fundamental task for NER, equipped with a classifier layer on top of the pre-trained BERT architecture. The use of BERT in this phase underscores our goal to capture the essence of the knowledge contained within the LLM-generated annotations and to utilize it in enhancing the model’s capability in a focused, resource-efficient manner.
Training with Original Data
We conducted a control experiment to establish a performance baseline for the BERT model. This was crucial for later comparison with the results obtained from the BERT model trained on the GPT-4 annotated data. For this purpose, we used another random subset of 1000 sentences from the CONLL2003 training set, excluding the portion annotated by GPT-4.
The intent behind using only the original dataset for training was to isolate the effect of the distilled data on the model’s learning capabilities. Our training parameters were meticulously set to optimize the learning process: the maximum sequence length was capped at 128 tokens, with a training batch size of 4, a validation batch size of 2, for a total of 10 epochs. The learning rate was set at 1e-05, with a max gradient norm of 10 to stabilize the training process.
Upon training, the model was evaluated on the CONLL2003 test set. The results showcased the inherent capability of the BERT model to capture and classify named entities, with an overall precision of 0.80, recall of 0.83, and an F1-score of 0.81. The evaluation scores across individual entity types—LOC, MISC, ORG, and PER—provided a granular view of the model’s performance, with the PER category achieving the highest F1-score of 0.92.
Training with Distilled Data Followed by Original Data
In refining our experiment’s second phase, we conducted a nuanced analysis of BERT’s performance when trained on a combination of distilled and original data. This phase was structured into two distinct training and evaluation steps to assess the efficacy of the distilled data from GPT-4 annotations in enhancing the model’s learning process.
Step 1: Initial Training with Distilled Data. BERT’s first training cycle utilized the 1000 sentences annotated by GPT-4. Despite employing consistent hyperparameters with the control experiment, the evaluation on the CONLL2003 test set resulted in a precision of 0.46, recall of 0.76, and an F1-score of 0.57. A closer look at the performance across different entity types revealed that the model achieved an F1-score of 0.76 for LOC, 0.14 for MISC, 0.65 for ORG, and 0.93 for PER, as depicted in the provided evaluation details. These scores indicated a substantial variance across categories, with particularly lower scores in the MISC category, suggesting that the distilled data alone may not provide enough variety for the model to generalize effectively across all entity types.
Step 2: Augmented Training with Original Data. Subsequent to the initial training, BERT was further trained on an additional 1000 sentences randomly sampled from the CONLL2003 training dataset. This dataset was identical to the one used in the control group to maintain consistency and ensure a meaningful comparison. After completing the second training cycle with the same hyperparameters and an additional 10 epochs, BERT’s performance on the test set improved significantly, achieving an overall precision of 0.83, recall of 0.87, and an F1-score of 0.85. The detailed scores for individual entity types were 0.88 for LOC, 0.60 for MISC, 0.84 for ORG, and 0.95 for PER, reflecting considerable improvements across the board and notably addressing the earlier shortcomings in the MISC category.
@@
The experimental results provided a compelling narrative on the effectiveness of using a combined dataset for training the BERT model (shown in Table 2). When the model was initially trained solely on the distilled dataset annotated by GPT-4, the overall test performance yielded an F1-score of 0.57, indicating that while the model could capture the essence of the data for certain entity types, the overall effectiveness was not optimal. Notably, the performance on individual entity types such as LOC and PER was relatively strong, with F1-scores of 0.76 and 0.93 respectively, suggesting that GPT-4’s annotations were particularly beneficial for these categories.
However, the MISC category presented a significant challenge, with an F1-score of only 0.14. This could be attributed to the model’s unfamiliarity with the diverse and less defined nature of the MISC entities, compounded by GPT-4’s inclination to generalize unfamiliar content under the MISC label, which adversely affected the overall metrics.
In contrast, when BERT was subsequently trained on a combination of distilled and original data, the model’s performance improved dramatically, achieving an F1-score of 0.85 in the final test evaluation. This improvement was consistent across all entity types, with even the MISC category showing a notable increase to an F1-score of 0.60. This two-step training process, using both GPT-4 annotated data and human-annotated data, proved to be more effective than using either dataset alone.
The results strongly suggest that leveraging distilled data from advanced LLMs like GPT-4, when complemented with original human annotations, can significantly enhance the performance of NLP models. This strategy effectively combines the broad contextual knowledge of LLMs with the nuanced understanding of human annotators, leading to a more robust and efficient named entity recognition capability. Such an approach offers a promising avenue for reducing costs and increasing efficiency in NLP applications, validating the integration of LLM insights into traditional NLP tasks.
Metric Pure CONLL Data Training Distilled Data Training Combined Data Training
Overall
   Precision 0.80 0.46 0.83
   Recall 0.83 0.76 0.87
   F1-Score 0.81 0.57 0.85

### v3

@@ Abstract
Emerging Large Language Models (LLMs) like GPT-4 have revolutionized Natural Language Processing (NLP), showing potential in traditional tasks such as Named Entity Recognition (NER). Our study explores a three-phase training strategy that harnesses GPT-4’s capabilities to enhance the BERT model’s performance on NER. Initially, GPT-4 annotates a subset of the CONLL2003 and additional BBC dataset without fine-tuning. We then train BERT using a mix of original and LLM-annotated data, analyzing the efficacy of LLM annotations against traditional methods.
The second phase involves comparative experiments with different training regimens, assessing the synergy between distilled and original data. We observe that sequential strategies, particularly a simple mix of training first with distilled data followed by original data, significantly boost performance. In the third phase, we investigate various data blending techniques, including sigmoid and power decay functions, to optimize the training process further. Our results indicate that a strategic mix of distilled and original data markedly elevates the NER capabilities of BERT.
Our approach presents a scalable methodology that reduces manual annotation costs and increases efficiency, making it especially pertinent in resource-limited and closed-network environments. The study concludes that while the ’Simple Mix’ strategy yields the best results, understanding its underlying mechanisms requires further research. Future work will also focus on refining prompt designs and enhancing annotation selection processes, aiming to extend our methodology to diverse NLP tasks.
@@ Introduction
Addressing these challenges, our research proposes a comprehensive solution utilizing knowledge distillation techniques. By leveraging LLMs for data annotation and incorporating these annotations in the training of smaller models like BERT, we enhance the performance of these models on traditional NLP tasks. Our approach mitigates the limitations of LLMs through the use of Chain of Thought (CoT) prompting to refine reasoning processes and generate more accurate annotations. Furthermore, the output format in the prompt for instructing the LLM is more efficient for generative model to predict nested or discontinues entities than merely outputting sequence of tags.
Through a series of meticulously designed experiments, we have undertaken a comprehensive evaluation of BERT model training strategies, leveraging both GPT-4 annotated (distilled) data and traditional human-annotated data. Our multifaceted approach included exploring the impact of various training regimens: training exclusively with distilled data, exclusively with original data, and a combination of both in varying sequences and proportions. We investigated the potential of distilled data from similar domains to enhance performance and employed innovative data blending techniques to optimize training efficacy. The results from these experiments have been illuminating, consistently demonstrating that strategic training with a mix of distilled and original data can significantly elevate the model’s NER capabilities. Notably, our findings suggest that simple sequential mixing of data types and fine-tuning the proportion of data sources throughout the training process can lead to substantial performance gains.
@@
The integration of Large Language Models (LLMs) [26, 21, 23, 27, 5, 4, 24, 25, 6, 20, 7, 9, 8, 15, 14, 8] into Named Entity Recognition (NER) tasks have seen varied approaches and methodologies, demonstrating the evolving landscape of NLP research in leveraging advanced AI models for enhancing traditional tasks. This section outlines related work that has contributed to the development of methods and strategies in applying LLMs and other advanced techniques to NER, highlighting the diverse efforts to address challenges such as data scarcity, model efficiency, and cross-lingual applications. Many LLMs further improve the multilingual natural language processing and understanding tasks by supervised fine-tuning [16, 12, 11, 20, 13].
Named Entity Recognition
@@
Cross-lingual NER remains a significant challenge due to the scarcity of annotated training data, especially in low-resource languages. To minimize the gap among different languages, some researchers [18, 2, al2015polyglot,yang2022crop, 3, 10, 19], introduce the translation-based framework to enhance zero-shot cross-lingual NER, demonstrating substantial improvements over previous approaches. The gap between the sequence labeling nature of NER tasks and the text-generation orientation of LLMs is explored in [28, 38, 30, 35, 34], where the GPT-NER model is proposed. By transforming the NER task into a text generation format, GPT-NER addresses the hallucination issue common in LLMs, showing promise in low-resource and few-shot settings. PromptNER [29] presents a novel few-shot and cross-domain NER algorithm that utilizes prompt-based heuristics with LLMs, requiring only a few-shot examples and entity type definitions to achieve state-of-the-art performance on several benchmarks. This underscores the potential of prompt-based methods in NER tasks. [36, 33] explores targeted distillation with mission-focused instruction tuning to train student models for NER tasks, demonstrating that these models can significantly outperform general instruction-tuned models across a broad application class, including open information extraction. The Multi-view Contrastive Learning approach for Cross-lingual NER [31, 37, 32, 38] tackles the challenge of semantic and token-level representation alignment across languages, achieving new state-of-the-art performance on the XTREME benchmark for 40 languages. Inspired by the Chain of Thought methodology, [39] leverages the step-by-step problem-solving capabilities of LLMs for Biomedical NER, incorporating external knowledge to enhance entity type determination, showcasing significant improvements in a domain-specific application. [41] focuses on zero-shot information extraction in NER, adapting reasoning strategies tailored for NER tasks with LLMs. The introduction of syntactic augmentation and a two-stage majority voting strategy demonstrates remarkable improvements across multiple benchmarks. Lastly, [42] addresses the issue of applying pre-trained transformer models to long documents for NER tasks. By generating a synthetic context retrieval training dataset and training a neural context retriever based on a BERT model, this work shows that it is possible to outperform unsupervised retrieval baselines, marking an advancement in handling long-document NER tasks. These works collectively underscore the dynamic interplay between LLMs, traditional NLP techniques, and innovative methodologies in advancing the field of NER. Our research builds upon these foundations, aiming to further leverage the capabilities of LLMs to enhance NER model training and performance, particularly through the integration of Chain of Thought prompting and knowledge distillation techniques.
@@
This research delineates a three-phase exploration into augmenting Named Entity Recognition (NER) performance by integrating Large Language Models (LLMs), specifically GPT-4, for data annotation, and subsequently training a BERT model with these annotations. Initially, we assessed GPT-4’s annotation capabilities on a subset of the CONLL2003 dataset[43], employing both standard and Chain of Thought (CoT) prompting techniques to measure the quality of LLM-generated annotations against human annotations. Following this, we conduct a comparative training experiment to evaluate the impact of LLM annotations on BERT’s NER capabilities, using a blend of original and LLM-annotated data. The study further innovates by incorporating external, similar domain data (from the BBC news dataset) to enrich the training material, thus exploring the potential of external datasets to enhance model performance. The third part of our experiment involves a examination of various training strategies to determine the most effective methodology for integrating distilled and original data. This approach aims to provide insights into the benefits and limitations of leveraging LLM annotations within traditional NLP frameworks. The data and code supporting the findings of this study have been made publicly available and can be accessed online [^1] https://huggingface.co/datasets/kobe1987/DLLM2TM.
@@
For our annotation process, we leveraged the GPT-4 turbo API to meticulously generate named entity annotations for the selected CONLL2003 data samples. This process involved employing two distinct prompting methods (detailed in subsequent sections and Figure 1 and Figure 2), which facilitated the model’s understanding and execution of the annotation task. While GPT-4’s outputs were predominantly accurate and aligned with the task requirements, a minimal amount of manual post-processing was necessary to rectify occasional irregularities in the generated annotations, ensuring the highest quality of data for our study.
@@
We employed a few-shot prompting approach, shown in Figure 1, with GPT-4 to annotate a random sample of 1000 sentences from the CONLL2003 dataset, serving as a baseline for comparison. Few-shot learning, a method where the model learns to perform tasks from a limited number of examples, was utilized to guide GPT-4 in accurately identifying and classifying named entities into four categories: LOC (Location), MISC (Miscellaneous), ORG (Organization), and PER (Person).
@@
Alongside the standard prompting approach, we implemented the Chain of Thought (CoT) prompting method combined with few-shot learning to annotate the same 1000 randomly sampled sentences from the CONLL2003 dataset using GPT-4, which is shown in Figure 2, . The CoT methodology is designed to mimic human-like reasoning by guiding the model through a stepwise thought process before arriving at a conclusion. This approach not only aims to improve the accuracy of the annotations by making the model’s thought process explicit but also seeks to provide insights into the model’s reasoning, thereby enhancing the interpretability of the annotations.
@@
Plus, as shown in both standard prompting and CoT prompting in Figure 1 and Figure 2, by utilizing a structured dictionary output format, LLMs can effectively capture and denote complex entity relationships within text. For instance, in the sentence "my fingers swelled up and hurt", LLMs can identify the discontinuous entity by outputting {’symptom’: ’fingers swelled up’, ’fingers hurt’}, encapsulating the related but non-adjacent terms. Similarly, for nested entities such as in "national basketball of America"(a made up organization name), the model can differentiate and simultaneously output the encompassing organization and the included location with {’ORG’: ’national basketball of America’, ’LOC’: ’America’}, demonstrating the prompt’s ability to recognize and represent nested entities accurately.
Figure 1: The Standard prompting Template
Figure 2: The CoT prompting Template
@@
In the next stage of our research, we utilized the BERT base model, specifically the uncased version known as bert-base-uncased, to fine-tune with GPT-4 annotated data. BERT (Bidirectional Encoder Representations from Transformers) is renowned in the field of NLP for setting new benchmarks across a variety of tasks. Its architecture incorporates 12 layers of bidirectional transformers, providing a nuanced understanding of contextual information within language.
The uncased model of BERT, pre-trained on an extensive corpus comprising the entire English Wikipedia and BookCorpus, utilizes self-supervised learning with two primary objectives: Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). These pre-training tasks allow BERT to develop a profound bidirectional representation of the English language, crucial for understanding the intricacies of natural language. BERT’s ability to process context effectively, coupled with its relatively compact architecture, makes it particularly suitable for scenarios with limited computational resources or in closed-network environments. In such cases, BERT can capitalize on the rich, context-aware insights extracted from LLMs like GPT-4 to enhance its performance significantly.
In the token classification task, the network we use harnesses the power of BERT’s transformer architecture. It begins with an embedding layer that combines word, position, and token type embeddings into a unified representation for each token. This is followed by a series of 12 transformer layers, each applying self-attention that dynamically weights the influence of different parts of the input sentence, refining the token representations based on the context provided by surrounding words.
The contextual representation of each token, h, is computed as a result of successive transformations applied by the encoder layers. Each token embedding is first passed through a series of self-attention heads within each transformer layer, which compute attention scores that reflect the token’s relationship with every other token in the sequence. The attention mechanism’s output for each token is a weighted sum of the value vectors, which is mathematically represented as:
$Attention(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$
where Q, K, and V are the query, key, and value matrices derived from the token embeddings, and $d_{k}$ is the dimensionality of the key vectors. The outputs from the attention heads are then concatenated and passed through a feed-forward neural network, then the token representations are normalized and passed up to the next transformer layer for further refinement. Upon completion of this process through all transformer layers, the final contextualized token representation h is obtained, which encapsulates both the semantic and syntactic information pertinent to each token within its sentence context. The classification layer then takes this representation and applies a linear transformation followed by a softmax function to produce a probability distribution over the possible entity tags:
[…]
where $W_{c}$ and $b_{c}$ are the weights and bias of the classifier layer, and h is the output from the final transformer layer.
Experiment Design
In the second phase of our comparative training experiments, we aimed to assess the effectiveness of using GPT-4 annotated data, termed as ’distilled data’, in conjunction with original data to enhance the BERT model’s performance on NER tasks.
Preparation: We began by sourcing additional data for annotation from the BBC news dataset [44], chosen for its topical similarity to the CONLL2003 dataset, which is also centered around news content and used to conducting research of document clustering in the original paper. This step was based on the premise that similar data types could bolster the model’s training when later combined with the original CONLL data. A pivotal observation from the previous phase indicated that longer sentences provided GPT-4 with the necessary context to produce more accurate annotations. Therefore, we specifically selected longer sentences for GPT-4 to annotate using the Chain of Thought (CoT) prompting method. After the annotation process and removing samples with formatting issues or lacking entities, we compiled a final set of 966 entries, distributed across categories: business, entertainment, politics, sport, and technology.
Experimental Setup: The experiment was structured into several groups, each representing a unique training regimen:
  • Group A utilized 1000 original sentences from the CONLL2003 dataset, training the BERT model for 20 epochs, to serve as a baseline representing training with human-annotated data only.
  • Group B involved training the model for 20 epochs solely on 1000 distilled sentences from the CONLL2003 dataset, providing insight into the utility of LLM-generated annotations.
  • Group C aimed to evaluate the impact of combining distilled data with data from a similar domain, training BERT for 20 epochs on 1000 distilled CONLL sentences augmented with 966 distilled BBC sentences.
  • Group D explored a sequential training approach, initially using 1000 distilled CONLL sentences for 10 epochs, followed by 1000 original CONLL sentences for a subsequent 10 epochs.
  • Group E followed a similar sequential pattern as Group D but started with the combined dataset of 1000 distilled CONLL sentences plus 966 distilled BBC sentences for the first 10 epochs, before proceeding with the 1000 original CONLL sentences for the remaining 10 epochs.
For each group, we employed a consistent set of hyperparameters to ensure comparability across results. The training and validation batches were set to sizes of 4 and 2, respectively, with a sequence length capped at 128 tokens. For each group, a fixed learning rate of 1e-05 and learning rate decay strategy with decreasing factor of 0.95 after each epoch are conducted respectively.
To address variability and ensure the robustness of our findings, each experimental setup was subjected to five iterations of training and testing. The average performance across these iterations was calculated and presented as the final result. The testing was conducted on the CONLL2003 test set, which comprises 3452 data points, allowing us to gauge the model’s NER capabilities after training under each experimental condition.
By conducting 50 distinct training cycles across various groupings and learning rate strategies, we sought to thoroughly investigate the potential synergies between distilled and original datasets in improving the BERT model’s NER performance.
@@
The analysis of the second phase’s experimental results (see Table 2) reveals several key insights into the performance of different training strategies on the BERT model for NER tasks.
Comparison of Pure Distilled Data Groups: Group B, trained only with distilled CONLL data, showed a micro average F1-score of 0.683 without learning rate decay, which is marginally lower than Group C’s score of 0.713 when trained on both CONLL and BBC distilled data without learning rate decay. This suggests that adding distilled data from a related external source (BBC) can enhance the model’s ability to generalize, as evidenced by the improvement in F1-scores.
Group C vs. Group A (Pure Original Data): Group C’s performance, even with the addition of BBC distilled data, did not surpass Group A, which trained purely on original data and achieved a micro average F1-score of 0.850 without learning rate decay. The variance in data sources may contribute to this outcome, emphasizing the importance of consistency.
Sequential Training Strategies: Groups D and E adopted a sequential training strategy, beginning with distilled data and then training on original data. Group D, which mixed CONLL distilled data with original data, showed an improvement in micro average F1-score to 0.859 without learning rate decay. Group E, which included BBC distilled data in the initial training phase, matched this performance, also achieving a micro average F1-score of 0.869 without learning rate decay. These results confirm that a sequential training approach can significantly elevate model performance compared to training solely on original data (Group A).
Learning Rate Decay Influence: The non-learning rate decay setup was generally more favorable for the sequential training groups. For Group E, many scores, e.g. the micro average F1-score and precision and the F1-score of LOC etc, are higher without decay. This pattern indicates that a consistent learning rate might be beneficial when transitioning from distilled to original data during training.
From these observations, it is evident that incorporating distilled data, particularly through a sequential training strategy, can yield significant improvements in NER performance. Moreover, the use of related external datasets as additional distilled data sources has the potential to further improve the model’s training efficacy.
Phase Three: Optimizing Training Strategy
Experiment Design
In Phase Three, we examine various strategies to integrate distilled data and original data throughout the training process. This phase aims to identify the most effective blending technique for enhancing the BERT model’s performance on NER tasks. To manage the proportions of data types throughout the training epochs, we introduce a function as follows:
$f(t)$ $=w_{0}$
$w_{1}$ $=1-w_{0}$
which determines the weight of distilled data used at any given epoch $t$. The weight of original data is then calculated as $w_{1}$, establishing a flexible framework that accommodates various blending strategies.
We explore seven distinct blending patterns, each defined by its characteristic function:
  • Pure Distilled Data(refer to Group C at phase two): For all epochs, the distilled data weight, $w_{0}$, remains at 1, signifying exclusive use of distilled data.
  • Pure Original Data(refer to Group A at phase two): Conversely, $w_{0}$ stays at 0, indicating the training uses only original data.
  • Simple Mix(refer to Group E at phase two): A direct transition is employed where $w_{0}$ is 1 for the first 10 epochs and 0 for the second 10 epochs, indicating an initial training with distilled data followed by training with original data. And the number of total epochs is represented with $T$.
    $w_{0}=\begin{cases}1&\text{if }t<\frac{T}{2}\\ 0&\text{otherwise}\end{cases}$
  • Sigmoid Decay: Here, $w_{0}$ follows a sigmoid function, smoothly transitioning from 1 to 0 as epochs progress, with a variable k determining the steepness of the curve. This pattern is represented by the equation:
    $w_{0}$ $=1-\frac{1}{1+e^{-k(x-0.5)}}$
    $x$ $=\frac{t}{T-1}$
    The experiment utilizes sigmoid functions with different $k$ values to observe the impact of transition steepness on training efficacy. Specifically, $k=\{2,4,8,16,32\}$ in our experiment.
  • Cosine Decay: Utilizing a cosine function, $w_{0}$ decreases in a sinusoidal pattern, providing a smooth and periodic blending of data. The function is defined as:
    $w_{0}=0.5\left(1+\cos\left(\pi\frac{t}{T-1}\right)\right)$
  • Power Decay: This method offers a variety of decay rates with the equation:
    $w_{0}=1-\left(\frac{t}{T-1}\right)^{n}$
    where different values of $n$ control the rate at which $w_{0}$ approaches 0. Specifically, $n=\{0.1,0.2,0.5,1,2,5,10\}$ in our experiment.
  • Full Blend (ALL): Not depicted in the graph, this approach completely mixes distilled and original data across all epochs, with $w_{0}=w_{1}=1$.
Each strategy’s parameters are informed by the figure (see Figure 3), visually illustrating the transformation patterns of $w_{0}$ across epochs for sigmoid, cosine, and power decay functions.
The training details for each pattern remain consistent with the previous phase, using the BERT model and the same hyperparameters, with a distilled dataset of 1966 sentences and an original dataset of 1000 CONLL training sentences. The learning rate decay is also applied as before.
Figure 3: Strategies to integrate distilled data and original data
Results and Analysis
The analysis of the third phase of our experiment, focusing on optimizing training strategies for BERT on NER tasks, reveals some intriguing findings. Here’s an in-depth look at the specific results presented in the table(see Figure 3):
Simple Mix Strategy Analysis: The ’Simple Mix’ approach without learning rate (LR) decay leads to the highest overall performance. Specifically, it achieves a micro-average F1-score of 0.869, a precision of 0.864, and a recall of 0.874. Notably, the LOC category benefits significantly from this strategy, with precision reaching 0.888 and recall at 0.913. However, this strategy does not always dominate across all metrics. For instance, the sigmoid approach with k=32 outperforms it on the macro-average recall with a score of 0.855 and the PER precision with a notable score of 0.957, suggesting that the sigmoid approach may offer competitive advantages in certain scenarios.
Learning Rate Decay Impact: Comparing the results with and without LR decay within the ’Simple Mix’ strategy, the non-decay approach generally yields better performance, as evidenced by the higher precision scores across most categories. This suggests that maintaining a constant learning rate might be more beneficial for model stability in certain contexts.
Sigmoid Strategy Evaluation: The sigmoid strategy with larger ’k’ values tends to perform better, approaching the ’Simple Mix’ results as k increases to 32. The sigmoid curve with k=32 is getting closer to the ’Simple Mix’ strategy in terms of performance, suggesting that a more gradual transition between the distilled and original data might provide a more effective learning trajectory.
Power Strategy Insights: For the power strategy, n=0.2 seems to offer the best overall results. This configuration leads to the highest scores in PER F1 (0.956), ORG F1 (0.849), and ORG recall (0.853). This might indicate that a moderately rapid shift towards the original data early in training benefits the model, possibly due to the distinct distributions of the original and distilled data.
Cosine Strategy Results: The cosine strategy provides moderate outcomes. The smooth transition offered by this strategy does not seem to offer any significant advantage over the others, indicating that it may not be the most effective for this application.
Comparison with ALL Strategy: The ’ALL’ strategy, which combines both datasets in full for each training epoch, shows the least effective performance. It suggests that simply mixing a larger volume of data from different distributions without a strategic approach can degrade model performance. This highlights the importance of a targeted training strategy over mere data quantity.
From these observations, we can conclude that while the ’Simple Mix’ strategy without LR decay stands out for its robust performance, particularly in the LOC and PER categories, alternative strategies like sigmoid (with a higher ’k’) and power (with n around 0.2) have their merits. The ’ALL’ strategy’s underperformance reinforces the need for strategic data blending rather than relying on data quantity alone. Moving forward, these insights will be invaluable for developing more nuanced training strategies that consider the nature and source of training data to maximize performance on NER tasks.
Group B
(pure CONLL distilled data)
Group C
(CONLL + BBC distilled data)
Group A
(pure original data)
Group D
(simple mix with CONLL only)
Group E
(simple mix with CONLL + BBC)
no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay
f1-score 0.683 0.672 0.713 0.691 0.850 0.849 0.859 0.853 0.869 0.869
precision 0.617 0.597 0.652 0.618 0.842 0.841 0.851 0.846 0.864 0.859
micro avg
(support: 8598)
recall 0.766 0.768 0.788 0.783 0.859 0.856 0.868 0.859 0.874 0.879
f1-score 0.660 0.655 0.684 0.674 0.820 0.818 0.829 0.820 0.840 0.839
precision 0.639 0.634 0.659 0.654 0.809 0.807 0.818 0.810 0.830 0.825
macro avg
(support: 8598)
recall 0.728 0.729 0.747 0.748 0.834 0.830 0.844 0.832 0.853 0.855
f1-score 0.735 0.732 0.756 0.750 0.851 0.850 0.861 0.854 0.871 0.871
precision 0.728 0.724 0.745 0.744 0.845 0.845 0.856 0.851 0.869 0.864
weighted avg
(support: 8598)
recall 0.766 0.768 0.788 0.783 0.859 0.856 0.868 0.859 0.874 0.879
f1-score 0.768 0.776 0.807 0.790 0.873 0.870 0.877 0.875 0.897 0.892
precision 0.711 0.719 0.762 0.745 0.856 0.851 0.862 0.854 0.888 0.872